# On Functional Dimension and Persistent Pseudodimension

### v1

@@ 1 Introduction
The upper bound above is a supremum over finite batches of points $Z=\{z_{1},\ldots,z_{m}\}$ of the row rank (over $\mathbb{R}$) of the Jacobian matrix associated to $Z,\theta$, viewed as a matrix over the polynomial ring $R$.
@@
The matrix $\alpha(\theta,Z)$ is called the activation matrix for $\theta,Z$. It was originally defined in [18] and incorporates information about the active paths in the computational graph for the architecture in a very natural way. See Sections 2.6 and 7 for more details.
These considerations, taken together, lead us to conjecture that the upper and lower bounds on the persistent pseudodimension of a parameter in a ReLU neural network class are equal, which would imply:
@@
$\mbox{dim}_{fun}(\theta)=dim_{p.{VC\Delta}}(\mathcal{F},\theta)$
@@
Fix a finite set $Z$ of points in input space $\mathbb{R}^{n_{0}}$. For any architecture of feedforward ReLU neural networks $\mathbb{R}^{n_{0}}\to\mathbb{R}$, locally near almost every parameter $\theta$, the parameter space $\Omega\cong\mathbb{R}^{D}$ is (diffeomorphic to) a product of a $\textrm{dim}_{\textrm{fun}}(\theta)$-dimensional submanifold consisting of parameters that all determine different functions on $Z$ and a $(D-\textrm{dim}_{\textrm{fun}}(\theta))$-dimensional consisting of directions in which perturbing the parameter does not change the function on $Z$.
A consequence of this theorem is that locally near a full measure set of parameters $\theta$, level sets of the empirical loss landscape with respect to the training set $Z$ are unions of these $(D-\textrm{dim}_{\textrm{fun}}(\theta))$-dimensional submanifolds, and gradient vectors for the loss function are constrained to lie in directions orthogonal to these submanifolds. The batch functional dimension can therefore be viewed as the local number of effective trainable parameters.
@@ Acknowledgments
Many thanks to Adam Klivans, Francois Malgouyres, Joachim Bona-Pellissier, and Matus Telgarsky for interesting conversations. JEG and KL were partially supported by NSF grant #2133822 through the SCALE MoDL program. JEG was partially supported by Simons grant #635578.
@@ 2.1 Preliminary definitions
The definition below gives several useful maps associated to any parameterized family, and the notation we will use for these maps. (In it, we use the notation $C^{0}(A,B)$ to denote the collection of continuous maps from a set $A$ to a set $B$.)
@@ 2.2 Functional dimension definitions
Denoting by $T_{{\theta}}(\Omega)\cong\mathbb{R}^{D}$ the tangent space to $\Omega$ at $\theta$, the rows of the Jacobian matrix ${\bf J}\mbox{E}_{Z}|_{\theta}$ are the gradient vectors with respect to the parameters of the output components of $\mbox{E}_{z_{i}}$ for $z_{i}\in Z$. Accordingly, they span a subspace of $T_{\theta}(\Omega)$, which we denote by $\mathcal{T}_{\theta}(Z)$. It is immediate from the definitions that $dim_{ba.fun}(\theta,Z)=\mbox{dim}(\mathcal{T}_{\theta}(Z))$.
@@ 2.4 Piecewise polynomial parameterized function families
In this subsection, we focus on parameterized families $\mathcal{F}$ that are not only continuous but also finitely piecewise polynomial (as is the case for ReLU neural network families). In this situation, since continuity implies that the pieces of $\Omega\times\mathbb{R}^{n_{0}}$ (which by definition contain a nonempty open subset) on which $\mathcal{F}$ is a polynomial are defined by polynomial equations; in fact, it is straightforward to show that the pieces are basic closed semialgebraic sets (by definition, a basic closed semialgebraic set is a set $S\subset\mathbb{R}^{n}$ given by finitely many simultaneous polynomial inequalities, $S=\{x\in\mathbb{R}^{n}\mid p_{1}(x)\geq 0,p_{2}(x)\geq 0,\ldots,p_{r}(x)\geq 0\}$ where $p_{i}(x)\in\mathbb{R}[x]$, $1\leq i\leq r$, are polynomials.) Note that i) both $\mathbb{R}^{n_{0}}$ and $\Omega$ are basic closed algebraic subsets of $\Omega\times\mathbb{R}^{n_{0}}$, the domain of $\mathcal{F}$, and ii) a finite intersection of basic closed semialgebraic sets is basic closed semialgebraic. Consequently, both every $F_{\theta}:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{d}}$ and every $E_{z}:\Omega\to\mathbb{R}^{n_{d}}$ are finitely piecewise-polynomial with pieces that are basic closed semialgebraic sets.
Since each piece of $\Omega\times\mathbb{R}^{n_{0}}$ contains an open set (by the definition of “piece”), the Identity Theorem for polynomials in several variables guarantees that there is a unique formal polynomial – an element of the ring $\mathbb{R}[\theta_{1},\ldots,\theta_{D},x_{1},\ldots,x_{n_{0}}]\,$ the polynomial ring over $D+n_{0}$ formal variables – that represents the restriction of $\mathcal{F}$ to the piece. In the definition below, we will use the following notation: given $r\in R$, $z\in\mathbb{R}^{n_{0}}$, $\tilde{\theta}\in\Omega$, we will denote by
@@
the element of the polynomial ring $R\coloneqq\mathbb{R}[\theta_{1},\ldots,\theta_{D}]$ obtained from $r$ by replacing the formal variables $x_{1},\ldots,x_{n}$ with the real numbers that are the corresponding coordinates of $z$; we will also denote by
@@
the element of the polynomial ring $\mathbb{R}[x_{1},\ldots,x_{n_{0}}]$ obtained by replacing the formal variables $\theta_{1},\ldots,\theta_{D}$ with the real numbers that are the corresponding coordinates of $\tilde{\theta}$. Henceforth, for convenience, we will use the notation
$R\coloneqq\mathbb{R}[\theta_{1},\ldots,\theta_{D}].$
@@
The reason for defining $E_{Z}^{R}$ by substituting in real numbers for formal variables in $P$ is that we want to keep track of how each variable appears in the polynomial, i.e. we need the map to be uniquely defined. (Because a polynomial in $D+n_{0}$ variables is not uniquely defined by specifying it’s values on a proper (e.g. $D$-dimensional or $n_{0}$-dimensional) linear subspace.)
@@ 2.5 ReLU networks: geometric and combinatorial background
which for almost all parameters is a hyperplane arrangement. $\mathcal{A}^{\ell}$ is said to be generic if for all subsets
@@
$\hat{\mathcal{A}}^{\ell}=\{\hat{H}^{\ell}_{i}\}_{i=1}^{n_{\ell}}:=\left\{F_{(\ell)}^{-1}(H_{i}^{\ell})\right\}_{i=1}^{n_{\ell}}\subseteq\mathbb{R}^{n_{0}}.$
@@
For $\ell\in\{1,\ldots,d\}$, denote by $R^{\ell}$ the polyhedral complex on $\mathbb{R}^{n_{\ell-1}}$ induced by the hyperplane arrangement associated to the $\ell^{\textrm{th}}$ layer map, $F^{\ell}$. Inductively define polyhedral complexes $\mathcal{C}\left(F_{(1)}\right),\ldots,\mathcal{C}\left(F_{(\ell)}\right)$ on $\mathbb{R}^{n_{0}}$ as follows: Set $\mathcal{C}\left(F_{(1)}=F^{1}\right):=R^{1}$ and for $i=2,\ldots,m$, set
$\mathcal{C}(F_{(\ell)}):=\left\{S\cap\left(F_{(\ell)}\right)^{-1}(Y)\mid S\in\mathcal{C}\left(F_{(\ell)}\right),Y\in R^{\ell}\right\}.$
@@ 2.6 Encoding algebraic representations via computational graphs
Following notation in Definition 4 of [22], let $F_{(\ell)}:=F^{\ell}\circ\ldots\circ F^{1}$ denote the composition of layer maps from the domain, ending with the $\ell$th layer map.
We refer to the components of $F_{(\ell)}$ as the neurons in the $\ell$th layer. The pre-activation map $y_{(\ell),i}:\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}$ associated to the $i$th neuron in the $\ell$th layer is given by:
@@ 2.7 Rank of matrices over polynomial rings
This is because two polynomials have some (iterated, partial) derivative that differs at a point if and only if they are different polynomials. (This consideration is also at the root of the reason why [14] defined batch functional dimension using only points $z$ that are parametrically smooth for a parameter $\theta$, as opposed to points $z$ so that $\mathcal{F}$ is once differentiable at $(z,\theta)$).
@@
$A\coloneqq\left\{\theta\in C\mid\textrm{rank}(E_{Z}(\theta))<r_{R}(\mathbf{J}E_{Z}^{R}|_{C})\right\}$
@@ 3.1.1 VC dimension and pseudodimension
$\Pi_{Z}(\mathcal{F}):=\{(\mbox{sgn}(F_{\theta}(z_{1})),\ldots,\mbox{sgn}(F_{\theta}(z_{m})))\,\,|\,\,\theta\not\in V_{Z}(\mathcal{F})\}$
@@
A finite set $Z=\{z_{i}\}_{i=1}^{m}\subset\mathbb{R}^{n_{0}}$ is shattered by $\mathcal{F}$ using the thresholds $t_{i}=F_{\theta}(z_{i})$ if and only if $Z$ is pseudo-shattered by the difference family $\mathcal{D}_{\theta_{0}}(\mathcal{F})$ at $\theta$.
@@ 3.1.2 Persistent pseudodimension
This issue also prevents a well-known proof technique for the Sauer-Shelah lemma (an induction argument based on an inclusion/exclusion principle) from applying directly to the persistent setting. Fortunately, the “shifting” proof of Sauer’s original result and discovered independently in [1, 11] does carry over to the persistent setting. The proof below sketches the shifting argument to verify that it carries over directly.

### v2

@@ 1 Introduction
Here, $E^{R}_{Z}(\theta)$ is the algebraic evaluation map (Def. 2.11), which outputs a vector whose $i$th entry is the polynomial expression for $F_{\theta}(z_{i})$ (as a function of $\theta$), and $\mathbf{J}$ denotes taking the Jacobian derivative of this map; thus, ${\bf J}E^{R}_{Z}(\theta))$ is a matrix whose $i$th row is the gradient with respect to the parameters. Applying $r_{\mathbb{R}}$ computes the real row rank (Def. 2.27) of this matrix, and the supremum is taken over all finite batches of points $Z=\{z_{1},\ldots,z_{m}\}$ in the domain.
@@
The matrix $\alpha(\theta,Z)$ is called the activation matrix for $\theta,Z$. It was originally defined in [30, 18] and incorporates information about the active paths in the computational graph for the architecture in a very natural way. See Sections 2.6 and 7 for more details.
These considerations, taken together, lead us to conjecture that the upper and lower bounds on the persistent pseudodimension of a parameter in a ReLU neural network class are equal.
@@
$\mbox{dim}_{fun}(\theta)=dim_{p.{VC\Delta}}(\mathcal{F},\theta).$
@@
Fix a finite set $Z$ of points in input space $\mathbb{R}^{n_{0}}$. For any architecture of feedforward ReLU neural networks $\mathbb{R}^{n_{0}}\to\mathbb{R}$, locally near almost every parameter $\theta$, the parameter space $\Omega\cong\mathbb{R}^{D}$ is (diffeomorphic to) a product of a $\textrm{dim}_{\textrm{ba.fun}}(\theta,Z)$-dimensional submanifold consisting of parameters that all determine different functions on $Z$ and a $(D-\textrm{dim}_{\textrm{ba.fun}}(\theta,Z))$-dimensional subspace consisting of directions in which perturbing the parameter does not change the function on $Z$.
(Here $\textrm{dim}_{\textrm{ba.fun}}$ denotes batch functional dimension, see Def. 2.7.) A consequence of this theorem is that locally near a full measure set of parameters $\theta$, level sets of the empirical loss landscape with respect to the training set $Z$ are unions of these $(D-\textrm{dim}_{\textrm{fun}}(\theta))$-dimensional submanifolds, and gradient vectors for the loss function are constrained to lie in directions orthogonal to these submanifolds. The batch functional dimension can therefore be viewed as the local number of effective trainable parameters for the batch.
We note that a special case of the theorem above is when the batch $Z$ is “big enough” and “well-chosen enough” that specifying the values of $F_{\theta}(z_{i})$ uniquely determines the function $F_{\theta}$ (at least locally, meaning that there is some open neighborhood $U$ of $\theta$ such that for any parameter $\theta^{\prime}\in U$, $F_{\theta^{\prime}}(z_{i})=F_{\theta}(z_{i})$ for all $\theta$ implies $\theta^{\prime}=\theta$). In this case, $\textrm{dim}_{\textrm{ba.fun}}(\theta,Z)=\textrm{dim}_{\textrm{fun}}(\theta)$, and we say that the functional dimension is attained on the batch $Z$. Proposition 5.16 of [14] proved that for a generic, transversal, combinatorially stable parameter $\theta$, functional dimension is attained on any decisive set $Z$, and gave an explicit definition of a decisive set for such a parameter.
@@ Acknowledgments
Many thanks to Adam Klivans, Francois Malgouyres, Joachim Bona-Pellissier, and Matus Telgarsky for interesting conversations. We also thank Francois Malgouyres for valuable comments on an earlier version of the paper. JEG and KL were partially supported by NSF grant #2133822 through the SCALE MoDL program. JEG was partially supported by Simons grant #635578.
@@ 2.1 Preliminary definitions
The definition below gives several useful maps associated to any parameterized family, and the notation we will use for these maps. (In it, we use the notation $C^{0}(A,B)$ to denote the collection of continuous maps from a topological space $A$ to a topological space $B$.)
@@ 2.2 Functional dimension definitions
Denoting by $T_{{\theta}}(\Omega)\cong\mathbb{R}^{D}$ the tangent space to $\Omega\cong\mathbb{R}^{D}$ at the point $\theta\in\Omega$, the rows of the Jacobian matrix ${\bf J}\mbox{E}_{Z}|_{\theta}$ are the gradient vectors with respect to the parameters of the output components of $\mbox{E}_{z_{i}}$ for $z_{i}\in Z$. Accordingly, they span a subspace of $T_{\theta}(\Omega)$, which we denote by $\mathcal{T}_{\theta}(Z)$. It is immediate from the definitions that $dim_{ba.fun}(\theta,Z)=\mbox{dim}(\mathcal{T}_{\theta}(Z))$.
@@ 2.3 ReLU neural network architectures as piecewise polynomial parameterized function classes
Given a composition of layer maps as in (1), we will (following [22, Definition 4]) use the notation
$F_{(\ell)}$ $\coloneqq F^{\ell}\circ\ldots\circ F^{1}.$ (3)
@@ 2.4 Piecewise polynomial parameterized function families
In this subsection, we focus on parameterized families $\mathcal{F}$ that are not only continuous but also finitely piecewise polynomial (as is the case for ReLU neural network families). A central theme of the results in this paper is proving relations (e.g. inequalities) between the ranks of various matrices whose entries are polynomial functions of the parameter $\theta$ and/or the input point $x\in\mathbb{R}^{n_{0}}$. Typically, the rank (whatever definition one is using of rank) of such a matrix will be constant for most values of $\theta$ and/or $x$, but there will certain values of $\theta$ and $x$ where the rank of the matrix will drop below this value (when, for example, specific values of $x$ and $\theta$ cause a linear dependency to develop between the rows or columns of the matrix). Thus, in order to prove results about the behavior of the parameterized function family in an open neighborhood of a point $(x,\theta)\in\mathbb{R}^{n_{0}}\times\Omega$, we may need to know not only the rank a matrix assumes at that specific $(x,\theta)$ (which could be one of the “bad” points where the rank drops), but the “typical” rank of the matrix on that neighborhood. To that end, we must investigate the polynomial expressions that describe the local behavior of the function family, and the matrices which have these formal polynomials as their entries, rather than their real-valued instances at specific points $(x,\theta)$. This becomes particularly important in Section 2.7.
In the case of finitely piecewise polynomial families, since continuity implies that the pieces of $\Omega\times\mathbb{R}^{n_{0}}$ (which by definition contain a nonempty open subset) on which $\mathcal{F}$ is a polynomial are defined by polynomial equations; in fact, it is straightforward to show that the pieces are basic closed semialgebraic sets (by definition, a basic closed semialgebraic set is a set $S\subset\mathbb{R}^{n}$ given by finitely many simultaneous polynomial inequalities, $S=\{x\in\mathbb{R}^{n}\mid p_{1}(x)\geq 0,p_{2}(x)\geq 0,\ldots,p_{r}(x)\geq 0\}$ where $p_{i}(x)\in\mathbb{R}[x]$, $1\leq i\leq r$, are polynomials.) Note that i) both $\mathbb{R}^{n_{0}}$ and $\Omega$ are basic closed algebraic subsets of $\Omega\times\mathbb{R}^{n_{0}}$, the domain of $\mathcal{F}$, and ii) a finite intersection of basic closed semialgebraic sets is basic closed semialgebraic. Consequently, both every $F_{\theta}:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{d}}$ and every $E_{z}:\Omega\to\mathbb{R}^{n_{d}}$ are finitely piecewise-polynomial with pieces that are basic closed semialgebraic sets.
Since each piece of $\Omega\times\mathbb{R}^{n_{0}}$ contains an open set (by the definition of “piece”), the Identity Theorem for polynomials in several variables guarantees that there is a unique formal polynomial – an element of the ring $\mathbb{R}[\theta_{1},\ldots,\theta_{D},x_{1},\ldots,x_{n_{0}}]\,$ the polynomial ring over $D+n_{0}$ formal variables – that represents the restriction of $\mathcal{F}$ to the piece.
Given a formal polynomial $r\in\mathbb{R}[\theta_{1},\ldots,\theta_{D},x_{1},\ldots,x_{n_{0}}]$, we will sometimes want to “plug in” real values for either the variables $\theta_{1},\ldots,\theta_{D}$ or the variables $x_{1},\ldots,x_{n_{0}}$; we now define notation for this. For
$r\in\mathbb{R}[\theta_{1},\ldots,\theta_{D},x_{1},\ldots,x_{n_{0}}],$
$z\in\mathbb{R}^{n_{0}}$ and $\tilde{\theta}\in\Omega$, we will denote by
@@
the element of the polynomial ring
$R\coloneqq\mathbb{R}[\theta_{1},\ldots,\theta_{D}]$
obtained from $r$ by replacing the formal variables $x_{1},\ldots,x_{n}$ with the real numbers that are the corresponding coordinates of $z$; we will also denote by
@@
the element of the polynomial ring $\mathbb{R}[x_{1},\ldots,x_{n_{0}}]$ obtained by replacing the formal variables $\theta_{1},\ldots,\theta_{D}$ with the real numbers that are the corresponding coordinates of $\tilde{\theta}$.
@@
The reason for defining the map $\theta\mapsto E_{Z}^{R}(\theta)$ by starting with a formal polynomial $P\in\mathbb{R}[\theta_{1},\ldots,\theta_{D},x_{1},\ldots,x_{n_{0}}]$ and then substituting in real numbers for formal variables in $P$ is that we want to keep track of how each variable appears in the polynomial, i.e. we need $E_{Z}^{R}(\theta)$ to be well-defined. The issue is that a polynomial in $D+n_{0}$ variables is not uniquely defined by specifying it’s values on a proper (e.g. $D$-dimensional or $n_{0}$-dimensional) linear subspace.
@@ 2.5 ReLU networks: geometric and combinatorial background
which for almost all parameters is a hyperplane arrangement (cf. [29]). $\mathcal{A}^{\ell}$ is said to be generic if for all subsets
@@
$\hat{\mathcal{A}}^{\ell}=\{\hat{H}^{\ell}_{i}\}_{i=1}^{n_{\ell}}:=\left\{F_{(\ell-1)}^{-1}(H_{i}^{\ell})\right\}_{i=1}^{n_{\ell}}\subseteq\mathbb{R}^{n_{0}}.$
(Recall that the notation $F_{(\ell)}$ is defined in (3).)
@@
For $\ell\in\{1,\ldots,d\}$, denote by $R^{\ell}$ the polyhedral complex on $\mathbb{R}^{n_{\ell-1}}$ induced by the hyperplane arrangement associated to the $\ell^{\textrm{th}}$ layer map, $F^{\ell}$. Inductively define polyhedral complexes $\mathcal{C}\left(F_{(1)}\right),\ldots,\mathcal{C}\left(F_{(d)}\right)$ on $\mathbb{R}^{n_{0}}$ as follows: Set $\mathcal{C}\left(F_{(1)}=F^{1}\right):=R^{1}$ and for $i=2,\ldots,m$, set
$\mathcal{C}(F_{(\ell)}):=\left\{S\cap F_{(\ell-1)}^{-1}(Y)\mid S\in\mathcal{C}\left(F_{(\ell-1)}\right),Y\in R^{\ell}\right\}.$
@@ 2.6 Encoding algebraic representations via computational graphs
Recall the notation $F_{(\ell)}:=F^{\ell}\circ\ldots\circ F^{1}$ from (3). We refer to the components of $F_{(\ell)}$ as the neurons in the $\ell$th layer. The pre-activation map $y_{(\ell),i}:\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}$ associated to the $i$th neuron in the $\ell$th layer is given by:
@@
(The notation $m(\gamma)$ was defined in the paragraph preceding that containing (7). See also Figure 1 for an example.)
@@ 2.7 Rank of matrices over polynomial rings
This is because two polynomials have (iterated, partial) derivatives that agree in a neighborhood of a point iff they are the same polynomial. (This consideration is also at the root of the reason why [14] defined batch functional dimension using only points $z$ that are parametrically smooth for a parameter $\theta$, as opposed to points $z$ so that $\mathcal{F}$ is once differentiable at $(z,\theta)$).
@@
$A\coloneqq\left\{\theta\in C\mid\textrm{rank}(\mathbf{J}E_{Z}(\theta))<r_{R}(\mathbf{J}E_{Z}^{R}|_{C})\right\}$
@@ 3.1.1 VC dimension and pseudodimension
$\Pi_{Z}(\mathcal{F}):=\left\{(\mbox{sgn}(F_{\theta}(z_{1})),\ldots,\mbox{sgn}(F_{\theta}(z_{m})))\,\,|\,\,\theta\not\in V_{Z}(\mathcal{F})\right\}.$
@@
A finite set $Z=\{z_{i}\}_{i=1}^{m}\subset\mathbb{R}^{n_{0}}$ is pseudo-shattered by $\mathcal{F}$ using the thresholds $t_{i}=F_{\theta}(z_{i})$ if and only if $Z$ is shattered by the difference family $\mathcal{D}_{\theta_{0}}(\mathcal{F})$ at $\theta$.
@@ 3.1.2 Persistent pseudodimension
This issue also prevents a well-known proof technique for the Sauer-Shelah lemma (an induction argument based on an inclusion/exclusion principle) from applying directly to the persistent setting. Fortunately, the “shifting” proof of Sauer’s original result, discovered independently in [1, 11], does carry over to the persistent setting. The proof below sketches the shifting argument to verify that it carries over directly.